# Personalized Subgraph Federated Learning with Differentiable Auxiliary Projections

### v1

@@ 4.3 Model Analysis
To intuitively show that the auxiliary projection vector APV can accurately capture the latent similarity among clients under non-IIDness, we construct a synthetic graph (See Sec. D.4) that jointly embodies the two types of heterogeneity: label heterogeneity and subgraph (structural and feature) heterogeneity. Results demonstrate that the APV-based client similarity can best recover the ground-truth client relationships. Each APV converges to the principal axis of its client’s embedding distribution, yielding highly aligned directions within the same group and near-orthogonal directions across different groups (Theorem 3.1). Thus APV can serve as a privacy-preserving yet information-rich descriptor in subgraph FL.
(e) Cora with 10 clients
(f) ogbn-arxiv with 10 clients
@@
To perform personalized FL, FED-PUB uses soft clustering by running a proxy graph through each client model and comparing embeddings. GCFL applies hard clustering, grouping clients with a Stoer-Wagner cut on cross-round gradients. Our APV-based method only computes similarity between the uploaded vectors, requiring no extra forward passes, gradient logging or cut computation, so server overhead is minimal. As Fig. 5 illustrates, the APV method attains the fastest clustering time and the lowest peak memory, making it more efficient at scale.
@@
Fig. 5 compares the convergence behaviour of FedAux and FedAvg. Notably, FedAux converges by the 60-th communication round in both the 10-client and 20-client settings. Since the latter involves a higher level of data heterogeneity, it indicates that FedAux maintains a consistent convergence speed even as the degree of non-IIDness increases.
In Appendix E, we conduct an ablation study to investigate the impact of our proposed continuous aggregation scheme, which encodes local node relationships in the 1D space induced by the APV, as well as the effect of the server-side APV-based personalized federated aggregation. We also conduct experiments to analyze the sensitivity to hyperparameters.
@@ 5 Related Work
For general Federated Learning, FedAvg (McMahan et al., 2017) first demonstrated that deep models can be trained on decentralized data with iterative model averaging. Subsequent work revealed that statistical and systems heterogeneity slow or destabilize FedAvg’s convergence, which has been formally addressed by proximal correction (Li et al., 2020), variance reduction (Karimireddy et al., 2020), and data–distribution smoothing through a small globally shared subset of samples (Zhao et al., 2018). Optimization refinements such as normalized aggregation (Wang et al., 2020), adaptive server updates (Reddi et al., 2021) and dynamic regularization (Acar et al., 2021) further tighten convergence guarantees under extreme non-IID settings. Beyond a single global model, personalisation frameworks, e.g., Ditto (Li et al., 2021), which jointly optimizes a shared model and client-specific objectives, explicitly trade off fairness, robustness, and local adaptation. For Graph Federated Learning (Xie et al., 2023; Li et al., 2024; Yue et al., 2024; Ye et al., 2023), at the node level, FedGCN (Yao et al., 2023) illustrates that one-shot encrypted exchange can suffice to federate GCNs while maintaining accuracy and privacy; GCFL (Xie et al., 2021) clusters clients by gradient dynamics to mitigate structural and feature shift across graphs. To combat cross-domain heterogeneity, FedStar (Tan et al., 2023) extracts a domain-invariant topology that generalizes across diverse graphs, and FedGraph (He et al., 2021) augments local data by requesting node information from other clients. When each participant owns only a fragment of a larger network, subgraph federated learning methods such as FedSage/FedSage+ (Zhang et al., 2021) generate virtual neighbours to repair missing cross-subgraph edges. FED-PUB (Baek et al., 2023) proposes to generate functional embeddings to evaluate the similarity between clients for personalized aggregation. These models collectively highlight an open challenge for personalized graph FL, i.e., accurate client similarity measures, which our proposed FedAux addresses through end-to-end learning of auxiliary projection vectors.
@@ Appendix A More Discussion of Related Work
GCFL (Xie et al., 2021) compares each client’s parameter updates before and after communication. If a client’s update deviates substantially from the majority, it is deemed too different and is excluded from the dominant aggregation cluster. However, GCFL relies on hard clustering, which can not capture finer-grained similarities across clients. Moreover, it depends on manually tuned hyperparameters to control how different a client must be before exclusion, leaving it unclear how large a deviation is tolerable without enough knowledge about the raw data. FED-PUB (Baek et al., 2023), on the other hand, constructs a proxy graph at the server and evaluates model outputs from each client on the proxy graph to measure their similarity. Clients with similar outputs on the proxy graph are deemed more alike and thus aggregated more closely. Yet building a suitable proxy graph as the middleware to measure the similarity between clients is nontrivial, because it requires simulating the real graph data on the server.
@@
In this work, we propose an orthogonal approach to measure and aggregate models on the server. Specifically, beyond distributing a global GNN (Zhuo & Tan, 2022b), the server also provides a global Auxiliary Projection Vector (APV) to each client at the start of every communication round, as shown in the left of Fig. 1. During local training, the APV is jointly optimized with the client’s GNN parameters to form a one-dimensional space onto which node embeddings are projected. This procedure tailors the APV, which adjusts the shape of this space, to the unique distribution of each subgraph by preserving its distinctive patterns. Once training concludes, clients upload their updated GNN parameters and local APV to the server, which then compares these learned APVs to gauge similarity and detect finer-grained, continuous heterogeneity across clients. By discarding the hard clustering thresholds used in GCFL and removing the need for a proxy dataset used in FED-PUB, our method offers a more flexible and efficient strategy to identify and aggregate similar subgraphs.
@@
Another advantage of our method lies in its privacy-preserving design. We argue that sharing detailed node embeddings, as in FedGCN (Yao et al., 2023), can leak private subgraph information, as adversaries on the server could compute pairwise similarities among embeddings to reconstruct local connectivity. Moreover, due to high dimensionality and lack of explicit structural encoding, directly comparing parameter matrices to measure raw data similarity is both unreliable and computationally unstable. In contrast, our proposed APV is a compact parameter vector that preserves essential node relationships without exposing the actual node embeddings, offering stronger privacy guarantees and structure-awareness. Additionally, its low-dimensional nature makes similarity computation more efficient and robust.
Appendix B Pseudo Code and Complexity Analysis for FedAux
The overall training algorithm is shown in Algorithm 1. For the client side of FedAux, the local GNN embedding generation incurs a complexity of $\mathcal{O}(|E_{k}|d^{\prime})$, and the auxiliary projection from embeddings to the APV results in $\mathcal{O}(N_{k}d^{\prime})$. Besides, the kernel-based embedding aggregation over the 1D space induced by the APV has complexity $\mathcal{O}(N_{k}^{2}d^{\prime})$. Consequently, the per-client complexity is $\mathcal{O}(|E_{k}|d^{\prime}+N_{k}^{2}d^{\prime})$. On the server side, computing the client-wise similarity for personalized federated aggregation involves a complexity $\mathcal{O}(K^{2}d^{\prime})$. Therefore, the total complexity of FedAux per communication round is $\mathcal{O}\left((|E_{k}|+N_{k}^{2}+K^{2})d^{\prime}\right)$.
@@ 5             $(\theta_{k},\bm{a}_{k})\leftarrow(\theta^{(t-1)}_{k},\bm{a}^{(t-1)}_{k})$;
7                   run local GNN forward pass with $h^{(t-1)}_{k,i}=f_{\theta^{(t-1)}_{k}}(v_{i})$ to learn node embeddings as Eq. 1;
@@
10                   compute kernel-based aggregated embeddings $z^{(t-1)}_{k,i}$ with Eq. 6;
11                   concatenate $\pi^{(t-1)}_{k,i}=[h^{(t-1)}_{k,i}\|z^{(t-1)}_{k,i}]$;
@@
      // Server-side personalised aggregation
20       compute $w^{(t-1)}_{k,l}$ with Eq. 7;
@@ Appendix C Proofs
C.1 Proof of Theorem 3.1
The proof relies on two mild structural assumptions, both typical in optimisation analyses of deep models.
@@
C.2 holds exactly for a linear-softmax classifier and is a first-order approximation for MLPs. Then we introduce the auxiliary lemmata.
@@
$s_{i}=\bm{a}^{\top}h_{i}$, with $\|\bm{a}\|=1$. Hence $\frac{\partial s_{i}}{\partial\bm{a}}=h_{i}$. Because we will always re-project $\bm{a}$ on the unit sphere after every update, the tangential component $h_{i}-s_{i}\bm{a}$ is the effective gradient, and the radial component vanishes. ∎
@@
Apply the chain rule to $\mathcal{K}_{ij}=\kappa(s_{i},s_{j})=\exp\left(-\frac{1}{\sigma^{2}}\left(s_{i}-s_{j}\right)^{2}\right)$ and invoke Lemma C.3 for $\partial(s_{i}-s_{j})/\partial\bm{a}$. ∎
Lemma C.5 (Gradient of the kernel-smoothed embedding).
@@
Let $D_{i}=\sum^{N}_{j=1}\mathcal{K}_{ij}h_{j}$, based on Eq. 6 we have $z_{i}=\frac{D_{i}}{M_{i}}$. Using the quotient rule, we have:
@@
Then we compute the two gradients in Eq. 15. The gradient of the $D_{i}$ w.r.t. $\bm{a}$ is $\frac{\partial D_{i}}{\partial\bm{a}}=\sum^{N}_{j=1}\frac{\partial\mathcal{K}_{ij}}{\partial\bm{a}}h_{j}^{\top}$, where $\frac{\partial\mathcal{K}_{ij}}{\partial\bm{a}}$ has been given by Lemma C.4. Thus we have:
@@
We can substitute Eq. 16 and Eq. 17 into the quotient rule Eq. 15:
@@
which is exactly the Eq. 14, completing the derivation. ∎
C.1.1 Proof of Theorem 3.1
@@
Using C.2 and Lemma C.5, we can perform chain-rule expansion of $\nabla_{\bm{a}}\mathcal{L}$ as:
@@
Substituting Eq. 14 into Eq. 21 gives:
@@
In Big-O notation, the symbol $\mathcal{O}\left(\sigma^{\mu}\right)$ as $\sigma\to 0^{+}$ means that there exists a constant $c>0$ and a neighbourhood $(0,\sigma_{0}]$ such that $|\mathcal{O}\left(\sigma^{\mu}\right)|\leq c\sigma^{\mu}$. Using C.1 and symmetry of the inner summation one obtains the compact matrix form of the above formula:
@@
with $\|\mathcal{R}(\sigma)\|=\mathcal{O}\left(\sigma^{0}\right)$. This proves Eq. 9.
Since $\bm{a}$ is re-normalised after every update, the effective tangential gradient (Delfour & Zolésio, 2011) is $\nabla_{\bm{a}}\mathcal{L}-(\bm{a}^{\top}\nabla_{\bm{a}}\mathcal{L})\bm{a}$. Note that $\bm{a}^{\top}\mathbf{C}\bm{a}$ is scalar, so subtracting the radial part yields the tangential gradient $-\mathbf{C}\bm{a}+\left(\bm{a}^{\top}\mathbf{C}\bm{a}\right)a$. A projected gradient descent step with learning rate $\eta$ therefore becomes:
@@
which is exactly the discrete-time Oja (Oja, 1982) update Eq. 10.
Let $\lambda_{\max}$ be the largest eigenvalue of $\mathbf{C}$, with unit-norm eigenvector $u_{\max}$. Standard theory of Oja’s algorithm (Krasulina, 1969; Oja, 1982) states:
  • Every eigenvector of $\mathbf{C}$ is a fixed point of Eq. 10.
  • All eigenvectors other than ±$\pm u_{\max}$ are unstable, and ±$\pm u_{\max}$ are globally asymptotically stable provided $0<\eta<2/\lambda_{\max}$.
Hence gradient descent drives $\bm{a}$ toward ±$\pm u_{\max}$. Because the kernel and the classifier do not depend on the sign of $\bm{a}$, both directions are equivalent, and choosing the positive-projection suffices. ∎
C.1.2 Interpretation of Theorem 3.1
Rayleigh-quotient maximization
Oja learning rule is a stochastic gradient ascent on the Rayleigh quotient $\mathcal{R}(\bm{a})=\bm{a}^{\top}\mathbf{C}\bm{a}$ over the unit sphere. Theorem 3.1 therefore formalizes the intuition that the APV aligns with the direction of maximum embedding variance.
@@
The leading term Eq. 10 is multiplied by $1/\sigma^{2}$. A smaller bandwidth increases the gradient magnitude, accelerating alignment but reducing smoothness; conversely, a larger $\sigma$ slows convergence while preserving differentiability.
@@
Once the APV converges to $u_{\max}$, the kernel weights $\mathcal{K}_{ij}$ depend only on $\left\langle h_{i},u_{\max}\right\rangle-\left\langle h_{j},u_{\max}\right\rangle$, which maximally separates nodes along the most informative one-dimensional projection. This precisely captures the fidelity property we desire.
C.2 Proof of Theorem 3.2
@@
By Lemma C.6 the non-diagonal weights vanish and $\mathcal{K}_{ii}\to 1$. Therefore $z_{i}\to h_{i}$. ∎
@@
Note that both matrices $\widetilde{\mathbf{Z}}$ and $\widetilde{\mathbf{H}}$ have the same ordering $\pi$ of rows. From Lemma C.7 each corresponding row converges as $\left\|z_{\pi(t)}-h_{\pi(t)}\right\|_{2}\rightarrow 0$ for every $t$. Since $N$ is finite, the Frobenius norm also converges to 0. ∎
C.2.1 Proof of Theorem 3.2
@@
The Conv1D operator $\mathrm{Conv}_{\mathbf{W}}$ is linear and its Lipschitz constant with respect to the Frobenius norm is $\textsc{Lip}_{\mathbf{W}}=\left(\sum_{\tau=1}^{B}\left\|\mathbf{W}_{\tau}\right\|_{2}^{2}\right)^{1/2}$, then for any two sequences $\mathcal{X}$ and $\mathcal{Y}$ we have:
@@
Lemma C.8 states that the right-hand side of Eq. 32 converges to 0, thus the left-hand side must converge to zero as well, establishing the claimed limit. ∎
C.3 Proof of Theorem 3.3

### v2

@@ 4.3 Model Analysis
To intuitively show that the auxiliary projection vector APV can accurately capture the latent similarity among clients under non-IIDness, we construct a synthetic graph (See Section˜D.4) that jointly embodies the two types of heterogeneity: label heterogeneity and subgraph (structural and feature) heterogeneity. Results in Fig.˜3 demonstrate that the APV-based client similarity can best recover the ground‑truth client relationships. Each APV converges to the principal axis of its client’s embedding distribution, yielding highly aligned directions within the same group and near‑orthogonal directions across different groups (Theorem˜3.1). Thus APV can serve as a privacy‑preserving yet information‑rich descriptor in subgraph FL.
(d) Cora with 10 clients
(e) ogbn-arxiv with 10 clients
@@
Membership Inference Attacks (MIA)
We add a comprehensive empirical privacy evaluation via MIAs to rigorously test whether our APV-based framework leaks sensitive information. We assume an honest‑but‑curious server with full access to the entire history of a client’s APV $\bm{a}_{k}$ and its GNN parameters $\theta_{k}$. The goal of the server is to decide whether a probe embedding $h_{\text{prob}}$ originated from client $G_{k}$, using only $(\bm{a}_{k},\theta_{k})$, while raw data, node embeddings, and gradients are never shared. Following the standard MIA methodology [21, 27], the server trains an attack classifier $g_{k}(\bm{a}_{k},\theta_{k},h_{\text{prob}})$ implemented as a two‑layer MLP with hidden size 64, to output the probability that $h_{\text{prob}}$ belongs to the training set of $G_{k}$. The attack is trained by maximum likelihood on a held‑out mixture of member against non‑member probes. In Table˜3, we compare FedAux against representative subgraph FL baselines on three datasets using identical client partitions to those in our main experiments, and report the Attack AUC (lower values indicate better privacy) averaged over five random seeds. The results show that MIAs against FedAux achieve AUCs in the range $[0.49,0.52]$, which is indistinguishable from random guessing. This demonstrates that APVs do not leak sensitive membership information and actually provide stronger privacy than baselines.
Transferability of APV
To examine whether APV can generalize beyond FedAux and serve as a plug-and-play personalization module, we integrate it into representative subgraph FL baselines. Specifically, we replace FED-PUB’s functional embedding similarity with APV, and jointly train APV with GraphGTA and FedSage+. All variants are trained under the same 10-client non-IID split on three datasets. Fig.˜4 shows that APV consistently improves each baseline across datasets. For example, GraphGTA with APV achieves 1%-2% higher accuracy than its original counterpart. These improvements confirm that APV functions as a generic, lightweight, and privacy-preserving similarity proxy that can be seamlessly integrated with diverse subgraph FL frameworks.
@@
To perform personalized FL, FED-PUB uses soft clustering by running a proxy graph through each client model and comparing embeddings. GCFL applies hard clustering, grouping clients with a Stoer-Wagner cut on cross-round gradients. Our APV-based method only computes similarity between the uploaded vectors, requiring no extra forward passes, gradient logging or cut computation, so server overhead is minimal. As Fig.˜6 illustrates, the APV method attains the fastest clustering time and the lowest peak memory, making it more efficient at scale.
@@
Fig.˜6 compares the convergence behavior of FedAux and FedAvg. Notably, FedAux converges by the 60th communication round in both the 10‑client and 20‑client settings. Since the latter involves a higher level of data heterogeneity, it indicates that FedAux maintains a consistent convergence speed even as the degree of non‑IIDness increases.
In Appendix˜E, we conduct an ablation study to investigate the impact of our proposed continuous aggregation scheme, which encodes local node relationships in the 1D space induced by the APV, as well as the effect of the server-side APV-based personalized federated aggregation. We also conduct experiments to analyze the sensitivity to hyperparameters.
@@ 5 Related Work
For general Federated Learning, FedAvg [20] first demonstrated that deep models can be trained on decentralized data with iterative model averaging. Subsequent work revealed that statistical and systems heterogeneity slow or destabilize FedAvg’s convergence, which has been formally addressed by proximal correction [14], variance reduction [11], and data–distribution smoothing through a small globally shared subset of samples [43]. Optimization refinements such as normalized aggregation [33], adaptive server updates [24] and dynamic regularization [1] further tighten convergence guarantees under extreme non-IID settings. Beyond a single global model, personalization frameworks, e.g., Ditto [15], which jointly optimizes a shared model and client-specific objectives, explicitly trade off fairness, robustness, and local adaptation. These advances establish the algorithmic and theoretical foundations on which federated graph learning is built.
For Graph Federated Learning [36, 17, 39, 38], at the node level, FedGCN [37] illustrates that one-shot encrypted exchange can suffice to federate GCNs while maintaining accuracy and privacy; GCFL [35] clusters clients by gradient dynamics to mitigate structural and feature shift across graphs. To combat cross-domain heterogeneity, FedStar [30] extracts a domain-invariant topology that generalizes across diverse graphs, and FedGraph [9] augments local data by requesting node information from other clients. When each participant owns only a fragment of a larger network, subgraph federated learning methods such as FedSage/FedSage+ [40] generate virtual neighbours to repair missing cross-subgraph edges. FED-PUB [3] proposes to generate functional embeddings to evaluate the similarity between clients for personalized aggregation. These models collectively highlight an open challenge for personalized graph FL, i.e., accurate client similarity measures, which our proposed FedAux addresses through end-to-end learning of auxiliary projection vectors.
Several new methods further extend personalization but come with non-trivial trade-offs. FedSSP [31] transmits spectral components that are invariant across domains but can leak sensitive local spectral information. FedEgo [41] shares ego-network embeddings, which directly expose structural patterns. PFGNAS [8] leverages prompt-based neural architecture search without server-side similarity estimation, but its reliance on LLM-based personalization introduces prohibitive cost. FedGrAINS [5] personalizes within each client by learning with a GFlowNet, thereby avoiding server-side mixing, but at the expense of training an additional model with high compute and memory overhead. While the abovementioned methods contribute interesting ideas, they either compromise privacy guarantees (e.g., FedSSP, FedEgo) or impose heavy computational costs (e.g., PFGNAS, FedGrAINS), limiting their practicality. In contrast, our method learns Auxiliary Projection Vectors (APV) in an end-to-end manner: lightweight, privacy-preserving client signatures that avoid raw data or embedding leakage, yet enable effective similarity-aware aggregation for personalized subgraph FL.
@@ 6 Conclusion
Acknowledgements
The research is supported, in part, by the Ministry of Education, Singapore, under its Academic Research Fund Tier 1 (RG101/24).
@@ Appendix A More Discussion of Related Work
GCFL [35] compares each client’s parameter updates before and after communication. If a client’s update deviates substantially from the majority, it is deemed too different and is excluded from the dominant aggregation cluster. However, GCFL relies on hard clustering, which can not capture finer-grained similarities across clients. Moreover, it depends on manually tuned hyperparameters to control how different a client must be before exclusion, leaving it unclear how large a deviation is tolerable without enough knowledge about the raw data. FED-PUB [3], on the other hand, constructs a proxy graph at the server and evaluates model outputs from each client on the proxy graph to measure their similarity. Clients with similar outputs on the proxy graph are deemed more alike and thus aggregated more closely. Yet building a suitable proxy graph as the middleware to measure the similarity between clients is nontrivial, because it requires simulating the real graph data on the server.
@@
In this work, we propose an orthogonal approach to measure and aggregate models on the server. Specifically, beyond distributing a global GNN, the server also provides a global Auxiliary Projection Vector (APV) to each client at the start of every communication round, as shown in the left of Fig.˜1. During local training, the APV is jointly optimized with the client’s GNN parameters to form a one-dimensional space onto which node embeddings are projected. This procedure tailors the APV, which adjusts the shape of this space, to the unique distribution of each subgraph by preserving its distinctive patterns. Once training concludes, clients upload their updated GNN parameters and local APV to the server, which then compares these learned APVs to gauge similarity and detect finer-grained, continuous heterogeneity across clients. By discarding the hard clustering thresholds used in GCFL and removing the need for a proxy dataset used in FED-PUB, our method offers a more flexible and efficient strategy to identify and aggregate similar subgraphs.
@@
Another advantage of our method lies in its privacy-preserving design. We argue that sharing detailed node embeddings, as in FedGCN [37], can leak private subgraph information, as adversaries on the server could compute pairwise similarities among embeddings to reconstruct local connectivity. Moreover, due to high dimensionality and lack of explicit structural encoding, directly comparing parameter matrices to measure raw data similarity is both unreliable and computationally unstable. In contrast, our proposed APV is a compact parameter vector that preserves essential node relationships without exposing the actual node embeddings, offering stronger privacy guarantees and structure-awareness. Additionally, its low-dimensional nature makes similarity computation more efficient and robust.
Appendix B Pseudo Code of FedAux
The overall training algorithm is shown in Algorithm˜1.
@@ 5    $(\theta_{k},\bm{a}_{k})\leftarrow(\theta^{(t-1)}_{k},\bm{a}^{(t-1)}_{k})$;
7         run local GNN forward pass with $h^{(t-1)}_{k,i}=f_{\theta^{(t-1)}_{k}}(v_{i})$ to learn node embeddings as Eq.˜1;
@@
10         compute kernel-based aggregated embeddings $z^{(t-1)}_{k,i}$ with Eq.˜6;
11         concatenate $r^{(t-1)}_{k,i}=[h^{(t-1)}_{k,i}\|z^{(t-1)}_{k,i}$];
@@
 // Server‑side personalised aggregation
20   compute $w^{(t-1)}_{k,l}$ with Eq.˜7;
@@ Appendix C Proofs
C.1 Proof of Theorem˜3.1
The proof relies on two mild structural assumptions, both typical in optimization analyses of deep models.
@@
˜C.2 holds exactly for a linear‑softmax classifier and is a first‑order approximation for MLPs. Then we introduce the auxiliary lemmata.
@@
$s_{i}=\bm{a}^{\top}h_{i}$, with $\|\bm{a}\|=1$. Hence $\frac{\partial s_{i}}{\partial\bm{a}}=h_{i}$. Because we will always re‑project $\bm{a}$ on the unit sphere after every update, the tangential component $h_{i}-s_{i}\bm{a}$ is the effective gradient, and the radial component vanishes. ∎
@@
Apply the chain rule to $\mathcal{K}_{ij}=\kappa(s_{i},s_{j})=\exp\left(-\frac{1}{\sigma^{2}}\left(s_{i}-s_{j}\right)^{2}\right)$ and invoke Lemma˜C.3 for $\partial(s_{i}-s_{j})/\partial\bm{a}$. ∎
Lemma C.5 (Gradient of the kernel‑smoothed embedding).
@@
Let $D_{i}=\sum^{N}_{j=1}\mathcal{K}_{ij}h_{j}$, based on Eq.˜6 we have $z_{i}=\frac{D_{i}}{M_{i}}$. Using the quotient rule, we have:
@@
Then we compute the two gradients in Eq.˜15. The gradient of the $D_{i}$ w.r.t. $\bm{a}$ is $\frac{\partial D_{i}}{\partial\bm{a}}=\sum^{N}_{j=1}\frac{\partial\mathcal{K}_{ij}}{\partial\bm{a}}h_{j}^{\top}$, where $\frac{\partial\mathcal{K}_{ij}}{\partial\bm{a}}$ has been given by Lemma˜C.4. Thus we have:
@@
We can substitute Eq.˜16 and Eq.˜17 into the quotient rule Eq.˜15:
@@
which is exactly the Eq.˜14, completing the derivation. ∎
C.1.1 Proof of Theorem˜3.1
@@
Using ˜C.2 and Lemma˜C.5, we can perform chain-rule expansion of $\nabla_{\bm{a}}\mathcal{L}$ as:
@@
Substituting Eq.˜14 into Eq.˜21 gives:
@@
In Big-O notation, the symbol $\mathcal{O}\left(\sigma^{\mu}\right)$ as $\sigma\to 0^{+}$ means that there exists a constant $c>0$ and a neighbourhood $(0,\sigma_{0}]$ such that $|\mathcal{O}\left(\sigma^{\mu}\right)|\leq c\sigma^{\mu}$. Using ˜C.1 and symmetry of the inner summation one obtains the compact matrix form of the above formula:
@@
with $\|\mathcal{R}(\sigma)\|=\mathcal{O}\left(\sigma^{0}\right)$. This proves Eq.˜9.
Since $\bm{a}$ is re‑normalised after every update, the effective tangential gradient [7] is $\nabla_{\bm{a}}\mathcal{L}-(\bm{a}^{\top}\nabla_{\bm{a}}\mathcal{L})\bm{a}$. Note that $\bm{a}^{\top}\mathbf{C}\bm{a}$ is scalar, so subtracting the radial part yields the tangential gradient $\mathbf{C}\bm{a}+\left(\bm{a}^{\top}\mathbf{C}\bm{a}\right)a$. A projected gradient descent step with learning rate $\eta$ therefore becomes:
@@
which is exactly the discrete‑time Oja [22] update Eq.˜10.
Let $\lambda_{\max}$ be the largest eigenvalue of $\mathbf{C}$, with unit-norm eigenvector $u_{\max}$. Standard theory of Oja’s algorithm [13, 22] states:
  • Every eigenvector of $\mathbf{C}$ is a fixed point of Eq.˜10.
  • All eigenvectors other than $\pm u_{\max}$ are unstable, and $\pm u_{\max}$ are globally asymptotically stable provided $0<\eta<2/\lambda_{\max}$.
Hence gradient descent drives $\bm{a}$ toward $\pm u_{\max}$. Because the kernel and the classifier do not depend on the sign of $\bm{a}$, both directions are equivalent, and choosing the positive‑projection suffices. ∎
C.1.2 Interpretation of Theorem˜3.1
Rayleigh‑quotient maximization
Oja learning rule is a stochastic gradient ascent on the Rayleigh quotient $\mathcal{R}(\bm{a})=\bm{a}^{\top}\mathbf{C}\bm{a}$ over the unit sphere. Theorem˜3.1 therefore formalizes the intuition that the APV aligns with the direction of maximum embedding variance.
@@
The leading term Eq.˜10 is multiplied by $1/\sigma^{2}$. A smaller bandwidth increases the gradient magnitude, accelerating alignment but reducing smoothness; conversely, a larger $\sigma$ slows convergence while preserving differentiability.
@@
Once the APV converges to $u_{\max}$, the kernel weights $\mathcal{K}_{ij}$ depend only on $\left\langle h_{i},u_{\max}\right\rangle-\left\langle h_{j},u_{\max}\right\rangle$, which maximally separates nodes along the most informative one‑dimensional projection. This precisely captures the fidelity property we desire.
C.2 Proof of Theorem˜3.2
@@
By Lemma˜C.6 the non-diagonal weights vanish and $\mathcal{K}_{ii}\to 1$. Therefore $z_{i}\to h_{i}$. ∎
@@
Note that both matrices $\widetilde{\mathbf{Z}}$ and $\widetilde{\mathbf{H}}$ have the same ordering $\pi$ of rows. From Lemma˜C.7 each corresponding row converges as $\left\|z_{\pi(t)}-h_{\pi(t)}\right\|_{2}\rightarrow 0$ for every $t$. Since $N$ is finite, the Frobenius norm also converges to 0. ∎
C.2.1 Proof of Theorem˜3.2
@@
The $\mathrm{Conv1D}$ operator $\mathrm{Conv}_{\mathbf{W}}$ is linear and its Lipschitz constant with respect to the Frobenius norm is $\textsc{Lip}_{\mathbf{W}}=\left(\sum_{\tau=1}^{B}\left\|\mathbf{W}_{\tau}\right\|_{2}^{2}\right)^{1/2}$, then for any two sequences $\mathcal{X}$ and $\mathcal{Y}$ we have:
@@
Lemma˜C.8 states that the right‑hand side of Eq.˜32 converges to 0, thus the left‑hand side must converge to zero as well, establishing the claimed limit. ∎
C.3 Proof of Theorem˜3.3